OpenStudy (anonymous):

need help again. man im not doing so well today.. f(x)= -x^2-9x-20.25 what are the x-intercepts. i have to show step by step

OpenStudy (anonymous):

Set it equal to zero and solve using complete the square or quadratic formula. If it makes it easier, you can divide it all by -1.

Parth (parthkohli):

You can rewrite the equation as: \(\Large \color{MidnightBlue}{\Rightarrow y = -x^2 - 9x - 20.25 }\) Get the y-intercept by setting x = 0 x-intercept by setting y = 0.

OpenStudy (anonymous):

QF again... becareful...:)

Parth (parthkohli):

Hmm?

OpenStudy (anonymous):

uh quick fox, quiz friday, quack funny, quadratic formula ... take your pick parth....

OpenStudy (anonymous):

Quarantined Foreigners

Parth (parthkohli):

\(\Large \color{MidnightBlue}{\Rightarrow y = x^2 + 9x + 20.25 }\) Since we have y = 0, \(\Large \color{MidnightBlue}{\Rightarrow x^2+ 9x + 20.25 = 0 }\)

OpenStudy (anonymous):

lol ^^^

Parth (parthkohli):

I like the formula :P

Parth (parthkohli):

\(\Large \color{MidnightBlue}{\Rightarrow x = {-b \pm \sqrt{b^2 - 4ac} \over 2a} }\)

OpenStudy (anonymous):

i did the quadratic formula and still didnt get the answers like in the book.

OpenStudy (anonymous):

i am bad at calculating

Parth (parthkohli):

You can do factoring then, if the formula is too tedious.

OpenStudy (anonymous):